The left and right page numbers of an open book are two consecutive integers whose sum is 209. Find these page numbers.


1
Expert's answer
2021-01-29T16:26:32-0500

Let "x" be the first page

2nd page ="x+1"


Thus ;

"x+x+1=209"


"2x+1=209"


"2x=208"


"x" "=104"

"x+1=105"


Thus the pages are 104 and 105